OpenStudy (anonymous):

Please help!?!???? Find an equation in standard form for the hyperbola with vertices at (0, ±6) and asymptotes at y = ± 3/4x?

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

\[ \frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1 \]

OpenStudy (anonymous):

h, k -> from vertices a,b from slopes of asymptotes

OpenStudy (anonymous):

(x-6)^2/9 + (y-k)/16= 1

OpenStudy (anonymous):

Did I do the denominators right? Please help me with the rest! I just don't get it

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

@Hero